OpenStudy (anonymous):

Solve the equation by finding the exact solution. ln x − ln 6 = 16

OpenStudy (anonymous):

three useful rules: \[\ln(a) - \ln(b) = \ln(a/b)\]\[\ln(a) + \ln(b) = \ln(ab)\]\[\ln(a^b) = b*\ln(a)\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and, if:\[c = \ln(d)\]\[d = e^c\]

OpenStudy (anonymous):

kinda lost me their

OpenStudy (jdoe0001):

notice the 1st rule listed by Euler271 ... that one will come very handy right off

OpenStudy (anonymous):

it would become ln(x/6)=16 correct?

OpenStudy (anonymous):

which will be x/6=e^16?

OpenStudy (anonymous):

it's just what ln(x) is: the inverse operation of e^x. solution: \[\ln(x) - \ln (16 )= 16\]\[\ln \left( \frac{ x }{ 6 } \right) = 16\]\[\frac{ x }{ 6 } = e^{16}\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

x = 6e^(16)

OpenStudy (anonymous):

sweet! thanks so much

OpenStudy (anonymous):

glad i could help:)

OpenStudy (anonymous):

your the best